A triangular-shaped wall with a base of 2x+4 & a height of x+3. The area of the triangle is 56 in.^2. What is the value of x?
Answer by amarjeeth123(569)

Area=1/2*base*height
1/2*(2x+4)*(x+3)=56
(x+2)(x+3)=56
x^2+5x+6=56
x^2+5x-50=0
x^2+10x-5x-50=0
x(x+10)-5(x+10)=0
(x-5)(x+10)=0
x=5 or x=-10
x=5 is the only possible solution.